OpenStudy (texaschic101):

I am not too good with these...did I do this right...can you work it out with proportions ??

OpenStudy (texaschic101):

in a spring the tension (T newtons) is directly proportional to its extension (xcm) When the tension is 150 newtons, the extension is 6cm (a) Find a formula for T in terms of x. T= ....? (B) calculate the tension, in newtons, when the extension is 15 cm newtons...? (C) calculate the extension in cm, when the tension is 600 newtons ....cm This is what I did... 150 newtons (T) = 6 cm (a) 150T = 6a T = 6a/150 B. 150/6 = x/15 cross multiply 6x = (150)(15) 6x = 2250 x = 375 newtons C. 150/6 = 600/x cross multiply 150x = (6)(600) 150x = 3600 x = 3600/150 x = 24 cm

OpenStudy (texaschic101):

@AccessDenied ....can you take a peek at this

OpenStudy (texaschic101):

@kropot72 ...is this correct ??

OpenStudy (texaschic101):

can this be solved with proportions or do I have to use the y = kx thing

OpenStudy (texaschic101):

@mathmale .....

OpenStudy (texaschic101):

I just noticed the first part is wrong....formula for T in terms of x....it can't be T = (6a)/150

OpenStudy (mathmale):

This is what I did... 150 newtons (T) = 6 cm (a) 150T = 6a T = 6a/150 Can you say what it is that you're looking for here? What's your purpose? You are given T and x (x being the extension); you are trying to find what quantity?

OpenStudy (texaschic101):

I do not know...I do not know how to do the y = kx thing ??

OpenStudy (texaschic101):

looking for the constant of proportionality ??

OpenStudy (texaschic101):

so is y newtons and the extension x ?

OpenStudy (mathmale):

that "y=kx thing" is exactly what you need. k is called "the constant of proportionality." Your job is to find k first, and then write a linear equation involving y, x and that k value. Try again: If T=150N and x=the spring extension= 6 cm., what is k?

OpenStudy (texaschic101):

y = kx 150 = 6k 150/6 = k 25 = k ??

OpenStudy (mathmale):

Yes, that's right. Now, with independent variable x and dependent variable T, write the required equation for T in terms of k and x.

OpenStudy (texaschic101):

T = 25x ?? Am I way off on this ??

OpenStudy (mathmale):

Not at all. Come on, show more confidence!

OpenStudy (texaschic101):

so that would be the answer for the first part ??

OpenStudy (texaschic101):

or would it be : x = T/25

OpenStudy (mathmale):

Look what you're trying to do: (B) calculate the tension, in newtons, when the extension is 15 cm newtons...? Using your new formula (answer for Part A), substitute x=15 cm and calculate T. Be certain to use the correct units of measurement.

OpenStudy (texaschic101):

T = 25x T = 25(15) T = 375 newtons thats what I got by doing a proportion. T = 25x 600 = 25x 600/25 = x 24 cm = x ok....I see now. Thank you so much mathmale :)

OpenStudy (mathmale):

My great pleasure! MM